# Foveated Retinotopy Improves Classification and Localization in CNNs

### v3

@@ Results
Fig 8: The visual search protocol is tested on the Animal 10k dataset (comprising $10,015$ images) and on the ImageNet validation dataset (comprising $50,000$ images). The mean accuracy of the networks is displayed, along with the accuracy in the pointing game.
Considering the localization task, we compute the likelihood map for each image and then select the fixation point with the maximum likelihood. The percentage of successful pointing games is calculated for each network, the two frameworks exhibit different performance profiles, with the retinotopic framework showing a pronounced capacity for efficient localization (Cartesian $78.74\%$, retinotopic $85.14\%$, see Figure 8). Furthermore, re-training the Cartesian network improves its localization accuracy (Cartesian $78.74\%$, “raw” $71.2\%$, see Figure 8).
We extend also this quantification on networks with different depths, ranging from ResNet $18$ to ResNet $101$. The trend observed for ResNet $101$ is also evident for shallower networks (see Table 1). The Cartesian networks show a slight advantage in categorization, while the retinotopic networks show improved performance in localization. It is noteworthy that smaller networks show improved localization capabilities, challenging the conventional wisdom that accuracy increases with the number of layers.
Table 1: Table with an analyses the key metrics for localisation (ratio activation and pointing game) and categorisation (accuracy before and after saccade with prior) as a function of the depth of the network.
@@
Before saccade 0.62 0.52 0.76 0.71 0.77 0.73
Saccade prior 0.92 0.86 0.95 0.94 0.96 0.95
@@
Exploring label association
@@
Fig 9: We tested the likelihood protocol on the Animal 10k (comprising $10,015$ images). (A) The mean likelihood across the point of fixation inside the bounding box (“In”) or the point of fixation outside the bounding box (“Out”) and the corresponding ratio of activation. (B) The intersection over union as a function of a threshold applied on the likelihood map.
Despite the increased diversity of poses, environments, and intra-class variability in Animal 10k compared to ImageNet, our results indicate that these methods effectively identify regions containing the target animals. A first observation is the activation of the likelihood map for this dataset. The peak threshold for the IoU is larger for the Cartesian network but both networks fine-tuned on the “focus” dataset express a better peak threshold (see Figure 9 (A)), although the Cartesian network obtains slightly better results on this metric. Also there the threshold is really close to $1$, expressing a more sustainable matching. When it comes to the contrast, as it seems in the average probability obtained inside (IN) versus outside (OUT) the areas of interest (see Figure 12), the retinotopic networks show the best contrast with the best ratio of activation (IN/OUT), also a similar dynamic for the networks fine-tuned on the “focus” dataset (see Figure 9-B).
Then, to demonstrate the accuracy of the pointing game for localization and the accuracy of the protocol, including saccades with a prior, for categorizing a label of interest. This allows the direct expression of both problems. All networks reach an accuracy of $1.00$ after a saccade to the position with the maximum likelihood value for the label “animal”. The highest improvement is found once within the retinotopic frame (“raw” $4.5\%$, Cartesian $4.1\%$, retinotopic $6.3\%$, see Figure 8). The pointing game results show that the retinotopic network outperforms the Cartesian network. With respect to these two metrics, the retinotopic framework shows higher localisation capabilities as observed in the ImageNet dataset. This suggests that our models have acquired general object-level features rather than dataset-specific cues. Testing on Animal 10k, with its greater pose diversity and difficulty, reinforces the conclusion that our method improves invariance to natural, ecologically valid geometric transformations beyond what standard CNNs alone can achieve.
@@
Fig 10: The likelihood maps presented here were generated by Resnet101 networks that were trained and tested on retinotopic space using the log-polar grid. (A) The original image, it is processed on a single iteration, with the likelihood maps displayed in accordance with the specific combinations of labels of interest. (B) Likelihood maps indicating the probability of the presence of the labels “dog”, “cat” and “bird” in the image. (C) The sample derived from all fixation points that yield the highest likelihood value for a label of interest.
@@
The likelihood map can be extended in the dimension of the label. In fact, the networks perform a discriminative categorization among $1000$ ImageNet labels. The likelihood maps generated by the model are expected to vary depending on the label or class predicted for a given input image. Specifically, the spatial distribution of high likelihood regions in the map differs depending on the label the model is trying to identify, or a class of objects defined by a set of labels. This suggests that the model learns to focus on different discriminative regions for different classes. The different likelihood maps provide insight into how the model’s spatial attention changes according to the visual patterns that drive each classification decision. We tested this hypothesis on an example image by showing likelihood maps for the classes “dog”, “cat”, and “bird”. As expected, the different regions are separated accordingly (see Figure 10).
@@ Discussion
The first and main result of this study is to demonstrate the excellent ability of standard deep CNNs to deal with foveated retinotopic inputs, even though this mapping enforces a radical transformation of the visual inputs. ResNet networks easily adapt to these inputs and the accuracy rates achieved with retinotopic inputs are equivalent to those of the original models. This is surprising given that the networks used in this re-training process were previously trained on Cartesian images and that images with a log-polar transformation show this high degree of distortion, in particular a high compression of visual information around the fixation point and a degradation of textures in the periphery, see Figure 1. One possible hypothesis is that the degradation of texture during the frame of reference change may cause the network to rely on shape rather than texture [44]. A further study could entail a comparison and replacement of the aforementioned parameters with those of humans, thereby providing insight into the evolutionary trajectory of these parameters.
@@
In addition, the log-polar transformation has the advantage of better invariance to zooms and rotations. This study shows first that for the original ResNet network on “regular” images, the average accuracy with a zoom or a rotation dropped sharply compared to baseline accuracy, confirming that this simple geometrical transformation mislead the networks. For rotations, the decline was steepest between two cardinal orientations, and demonstrate limited rotational invariance compared to humans [45, 46]. The robustness differs slightly when studying the impact of zooms, and this may be attributed to the transformation under examination that already provides a zooming-in effect of the fovea which accentuates the information in the area surrounding the point of fixation. To test this hypothesis, we may train a retinotopic version without applying the logarithmic function to the eccentricity axis. Finally, this invariance comes at the cost of reduced invariance to translations. For images that are not centered on the region of interest, one would have to shift the fixation point to the region of interest, similar to eye saccades. The initial consequence of this transformation can be observed in a comparison between the “regular” version of the networks and the “focus” version.
The integration of a retinotopic mapping approach holds great promise for improving the efficiency and accuracy of existing image processing tasks. First, it involves only a simple preprocessing of the input images, transforming them into novel images, and is therefore potentially applicable to existing image processing algorithms. The log-polar compression used in our approach allows a seamless extension to larger images without a significant increase in computational cost. On the other hand, this biological inspiration results in a more frugal energy consumption. This is an essential aspect when porting such algorithms to embedded systems under high power constraints. Overall, our results are consistent with features of biological vision. In particular, it is consistent with the robustness observed in physiological data on ultrafast image categorization, which show that human accuracy in recognizing briefly flashed images of animals is robust to rotation [46, 47].
@@
The second result of this study is the emergence of localization properties in networks re-trained with foveated inputs. Thanks to the definition of likelihood maps, that collates the output of the model while scanning the visual scene at a limited number of fixation points, this allows us to gain insight into the specificities of retinotopic processing: this transformation provides a more focused view, thus better separating the different elements of the image when focusing on its specific parts. This is reminiscent of pre-attentive mechanisms which allows biological vision to only porocess the important zones of the visual space. As a result, it provides a proxy for the measurement of saliency, in particular with respect to a set of cued labels.
@@
One hyopthesis is that the foveated retinotopic mapping implemented in the log-polar transform implements an efficient prior for visual object geometry. Indeed, a main source of variability in the view of an object comes from displacements of the observer relative to it, for example as the objects scales in the visual field as one approaches it, or as the object rotates with a rotation of the head. The log-polar mapping allows for a more invariant representation which explictly implements this prior for these displacements. Additionally, imposing translation invariance of the representation implements a prior on the possible representations allowed by the network [41]. However, it was observed that some classes have different statistics, and the relative size of buses are on average bigger than that of cats. Globally, we expect that different foveated retinotopies may emerge in different ecological niches.
@@
In our case, it seems that the foveated retinotopic mapping allows for a more precise localization of the category of interest, compared to off-the-shelf pre-trained networks using a Cartesian representation. It also gives us insight into the features on which our networks actually rely. Such information can be compared with physiological data [48], and be used to design better CNNs, and ultimately allow physiological tests to be proposed to further explore the features needed to classify a label of interest. In particular, by focusing on the point of fixation with the highest probability in the likelihood maps instead of the center of the bounding box defined in the “focus” dataset, we could consider refining the training of the network to our retinotopic mapping in a semi-supervised fashion.
Introducing eye movements to AI
@@
Building on these observations, simulating human saccadic eye movement patterns during visual tasks provides an exciting opportunity to gain further insights into these mechanisms using such networks. A protocol that iteratively classifies image patches corresponding to foveated regions in a manner that mimics eye movements could reveal how network performance is spatially modulated across the visual field. Comparing classification accuracy under different saccadic planning strategies, such as selecting the most uncertain or most likely location at each step, would provide valuable information about how network attention operates. This framework also allows the modeling of popular biological strategies for producing saccades, allowing a direct comparison with humans’ visual search behavior.
@@
Overall, implementing foveated classification with algorithmic saccades would provide a powerful method for validating existing attentional mechanisms in these networks, as well as inspiring new architectural innovations through embodied, task-driven visual attention modeling. In particular, this line of research should be particularly well fitted for a dual-pathway model which is well adapted to infer ego-motions [49, 50]. Finally, the implementation of this robust categorization, coupled with refined localization of a label of interest and optimal saccade selection, could allow us to extend this study to a more complex task. One such task is visual search (i.e., the simultaneous localization and detection of a visual target), and the likelihood maps could provide the underlying pre-attentive mechanisms on which its effectiveness seems to depend.
@@
In most mammals and amphibians, the arrangment of the visual field is preserved in the early visual pathway, a feature called retinotopy. Retinotopic mapping results from the combined effect of the arrangement of photoreceptors in the retina and their output convergence via the optic nerve. This causes nearby regions of the visual field to activate adjacent neural structures as signals travel from the retina to the brain. These mappings differ from species to species, and our study concentrates on foveated vision (as in humans) which gives more resolution to the central field of view. In particular, we implement it by transforming the Cartesian pixel coordinates into log-polar coordinates [27].
This simple parameterized transformation models accurately this biologically inspired retinotopic mapping. Considering arbitrary images (potentially with multiple channels such as RGB), each pixel’s position is defined by two real coordinates $(x,y)$ on a Cartesian reference frame. By convention, $x$ and $y$ are here considered belonging to the interval $[-1,1]$, with $(0,0)$ being the center of the image. To implement the concentration of pixels near the center of the retina, we need to consider an irregular grid in the Cartesian referential that maps to a regular grid in the log-polar referential. In the log-polar referential, the location of each pixel has corresponding coordinates (${\rho},{\theta}$) defined by:
@@
with $(x_{0},y_{0})$ defining the “center of fixation” (see Figure 1[51]. Importantly, $\rho$ and $\theta$ are only defined for $(x,y)\neq(x_{0},y_{0})$. In most of our experiments, we consider $x_{0}=(0,0)$, allowing us to focus on the central part of the image. Each dot ($\rho$, $\theta$) in the log-polar coordinate system has thus a unique correspondence in the Cartesian coordinate system (and vice versa): For each ($\log\rho$, $\theta$) belonging to the grid, the corresponding pixel coordinate is $(x_{0}+\rho\cos\theta,y_{0}+\rho\sin\theta)$.
In practice, images have a finite resolution and to avoid biases in the evaluation between networks, the number of angles sampled ($N_{\theta}$) and the number of eccentricities sampled ($N_{\rho}$) are set to $224$, so that the size of the transformed image match the resolution $224\times 224$ of the input images. This transformation is performed with the PyTorch library [52] through the use of the grid_sample() function, which maps the pixels of an input image to the coordinates of any arbitrary grid, using a linear interpolation to estimate the value of the pixels. This function is used, for instance, in spatial transformer networks [53].
Let’s now define each coordinates. All $\theta$ values are within a linear distribution in $[0;2\pi)$, while $\rho$ values are within a logarithmic interval with $r_{\text{min}}$ the minimal radius and $r_{\text{max}}$ be the maximal radius (with $N_{\rho}$ the radial resolution). In practice, we use a log-polar grid with an outer log-radius of $\log_{2}r_{\text{max}}=0$ ($r_{\text{max}}=1$, defining a circle tangent to the image box) and an inner log-radius of $\log_{2}r_{\text{min}}=-5$ ($r_{\text{min}}=2^{-5}$). In summary, the regular grid is the interval $[\log_{2}r_{\text{min}},...,\log_{2}(r_{\text{min}}+i\times\frac{r_{\text{max}}-r_{\text{min}}}{N_{\rho}-1}),...,\log_{2}r_{\text{max}}]$, for $i$ in $[0,...,N_{\rho}-1]$, in the $\log_{2}\rho$ dimension and $[0,...,j\times\frac{2\pi}{N_{\theta}},...,2\pi(1-\frac{1}{N_{\theta}})]$, for $j$ in $[0,...,N_{\theta}-1]$ in the $\theta$ dimension. Note that some pixels in the log-polar grid may be smaller than the pixels from the Cartesian grid.
@@
Convolutional neural networks (CNNs) have become essential tools in image classification, with several pre-trained models available for download. For example, the VGG family, including VGG16 and VGG19, introduced by Simonyan [54], uses deep architectures with $16$ or $19$ layers, consisting of stacked convolutional layers followed by fully connected layers. ResNet (Residual Networks), introduced by He et al. [33], addresses the vanishing gradient problem in deep networks by incorporating skip connections, allowing for the training of extremely deep networks (e.g. ResNet $50$, ResNet $101$ with respectively $50$ and $101$ layers). These models are widely used due to their robustness and scalability. The implementation of these deeper networks have demonstrated that deeper networks display enhanced resilience; however, this improvement is coupled with an overall increase in computational complexity [33]. Therefore, based on these findings, we focus on the deep CNN ResNet (with $18$ to $101$ layers from the PyTorch library  [^1] We further introduced a circular padding in the convolutions, however we controlled this had little impact overall (not shown).) on the ImageNet [32] categorization challenge which consists in classifying natural images into $1000$ labels.
@@
Two datasets were used for our study: the first dataset is the one from the ImageNet [32] challenge, which is the most widely benchmarked due to its huge collection of images and associated labels (the subset of ImageNet used in this study, i.e. ImageNet-1K with $1000$ labels and about $1000$ examples per label). In addition to the classification task, we consider here the localization task, which is prominent in computer vision. It consists in identifying both the label and the position of one or more objects of interest in an image. A distinction can be made between (i) the visual search task, where the label is given in advance and the task is simply to find the position of the object in the image, and (ii) the image labeling task, which consists in identifying both the objects and their position in the image, in the form of a labeled bounding box (see Figure 11). Despite its advantages, ImageNet has limitations for localization tasks. For example, the dataset lacks multi-labeling, with only one label per image. It is worth mentioning that ImageNet has some biases, the objects of interest are generally centered in the images, and the proportion of bounding boxes relative to the image size is relatively large, which may limit the impact of certain analyses.

### v4

@@ Results
Figure 8: The visual search protocol is tested on the ImageNet-1K validation dataset (comprising $50,000$ images) (left) and on the Animal 10k dataset (comprising $10,015$ images) (right). The mean accuracy of the networks is displayed, along with the accuracy in the pointing game.
Pointing game
We have shown that the likelihood maps provide us with an indication, for each image, of the best fixation point to identify the object we are looking for. However, this spatial indication (where to place the eye) does not tell us anything about the visual content of this fixation point. To estimate this visual content more precisely, we now consider the “pointing game” metric, that is the rate of successful pointing (i.e. landing inside the bounding box) when choosing the highest likelihood position on the likelihood map. The bounding boxes provide us with rather coarse information about the visual content of the image, dividing it into two zones: a box within which the object is present, and a box outline where it is assumed absent. This binary information (“in” or “out”) is, however, sufficient to indicate whether the point of highest likelihood is correctly located on or in the immediate vicinity of the object of interest.
We compute for each network on the validation set the percentage of successful pointings. The results are shown on Figure 8 (green bars, left panel), on the raw, Cartesian and Retinotopic networks. The “raw” network has $71.2\%$ success, the Cartesian has $78.74\%$ success and the Retinotopic $85.14\%$ success. Those rather high pointing game values reflect the fact that, on average, the bounding boxes cover a large portion of the image in the Imagenet-1k dataset. Nevertheless, the retinotopic network shows pointing scores significantly higher than those of the raw network and the Cartesian network. The position of maximal likelihood appears to be strongly correlated with the presence of the object, and this effect is notably sharper and more precise in the case of the retinotopic network.
We extend this visual search strategy by considering the classification accuracy obtained when placing the eye at the position of highest likelihood on the likelihood map, interpreted as a “saccade” toward the searched object. By doing so, the accuracy of all networks increases to about $96\%$ (“raw” $97.7\%$, Cartesian $96.2\%$, retinotopic $95.8\%$, see Figure 8), providing an estimation of the improvement that can be expected, with respect to the baseline, when placing the eye (or camera) at its highest likelihood position (interpreted as the highest likely position). Consistently with our previous results, the accuracy improvement is higher in the retinotopic case, with $22.9\%$ improvement, to be compared with $18.7\%$ and $18.9\%$ improvement in the “raw” and Cartesian cases respectively (see Figure 8). This strong response improvement reflects the critical role of the “direction of sight” in image classification, even in the Cartesian case, an issue little evoked in mainstream computer vision. It also suggest that the retinotopic network could be further improved by introducing such saccades during the retraining phase.
Impact of network depth on localization and classification
Table 1: Analyses of the key metrics for localisation (in/out likelihood ratio and pointing game) and classification (central fixation and max likelihood (“saccade”) fixation) for different network depths.
@@
Central fixation 0.62 0.52 0.76 0.71 0.77 0.73
With saccade 0.92 0.86 0.95 0.94 0.96 0.95
@@
We extend the analysis by considering network depths, ranging from ResNet $18$ to ResNet $101$. The results are shown in Table 1. We observe the same general trend across different network sizes: classification accuracy improves when the fixation point is placed at the position of highest likelihood, likelihood values are higher inside the bounding box, and the fixation point is predominantly located within the bounding box. These trends are present in both types of networks (Cartesian or Retinotopic), but they are more pronounced in the retinotopic networks. Network size primarily affects the classification rate, with lower rates observed for smaller networks. Classification rates remain slightly higher for Cartesian networks in all cases, but the improvement in classification rate based on fixation position is consistently stronger for retinotopic networks. Regarding localization (likelihood ratio and pointing game), network size does not appear to be a determining factor, with higher pointing scores and likelihood ratios observed for smaller networks, and again, a systematic advantage for retinotopic networks. In summary, Cartesian networks maintain a slight advantage in categorization, while retinotopic networks show improved performance in localization. It is noteworthy that smaller networks demonstrate better localization capabilities, challenging the conventional wisdom that performance necessarily increases with the number of layers.
@@
Figure 9: We tested the visual search protocol on the Animal 10k (comprising $10,015$ images). (A) The mean likelihood across the point of fixation inside the bounding box (“In”) or the point of fixation outside the bounding box (“Out”) and the corresponding likelihood ratio. (B) The intersection over union as a function of a threshold applied on the likelihood map.
Beyond ImageNet, the Animal-10k dataset
ImageNet-1K provides rich semantic links that allow the construction of task-specific datasets. It has previously been demonstrated that the use of fine-tuning to re-train networks such as Vgg16 (Simonyan and Zisserman, 2015) allows them to be applied to different tasks using the semantic network underlying ImageNet’s labels. Furthermore, it has been shown that the probability of a trained network performing a novel task (such as categorizing an animal) can be predicted using this semantic network, which links the outputs of a pre-trained network to a library of labels. This approach has been shown to be an effective method for learning to predict the presence of animals in images (Jérémie and Perrinet, 2023). We can then exploit the semantic connection underlying the ImageNet labels by using these networks to perform a categorization task on the second dataset, the Animal 10k dataset, as referenced by Yu et al. (2021), with the goal of identifying the animals depicted in the images. This dataset contains $10,015$ images, each depicting an animal. All images were integrated in a validation dataset to test the networks. The output of each network is interpreted as “animal” if the label represented by the highest likelihood after softmax is indeed an animal. We use this prediction to calculate the accuracy of predicting an animal for all networks. Note that the accuracy is slightly biased due to the absence of distractors, it only represents the probability of finding an animal when there is actually an animal in the input. Figure 8 (right panel) compares the accuracy improvement for both networks. The general accuracy is high (around $95\%$ in all cases) due to our simpler binary (animal/non-animal) classification task. Saccade improvement is still significant, with all networks reaching a maximal accuracy of $100\%$ after a saccade to the maximum likelihood position for the label “animal”. The highest improvement is once again found in the retinotopic case (“raw” $4.5\%$, Cartesian $4.1\%$, retinotopic $6.3\%$).
Regarding the localization capability, the In/Out likelihood ratio (Figure 9A) shows clear contrast enhancement, from 1.2 (raw network) toward 1.5 (Cartesian “focus” network) and 1.8 (Retinotopic “focus” network). The general lower contrast, when compared to ImageNet, comes again from the simpler binary (animal/non-animal) classification task, making it possible to guess the label “Animal” outside the bounding box from the background information. Looking now at the IoU metric (Figure 9B), a different trend is observed, with monotonically increasing IoU with the likelihood map threshold, reflecting a convergent matching of the likelihood map with the animal contour. The optimal threshold, close to 1, reflects a clear and sharp transition in both cases. Here, the retinotopic network provides slightly higher IoU values, except at the highest threshold case where the Cartesian network prevails. The difference between this and ImageNet lies in the fact that masks are used in the calculation instead of boxes (see Methods). Last, the pointing game results (see Figure 8B, green bars) provide a clear assessment of the Cartesian/retinotopic contrast. While both the raw and Cartesian network remain quite low (73-74% success), the retinotopic framework shows much higher pointing success (87.7%) despite refined animal contour, once again suggesting a more reliable animal localization. Despite the increased diversity of poses, environments, and intra-class variability in Animal 10k compared to ImageNet-1K, our results indicate that the regions containing the target animals are effectively identified.
@@
Figure 10: The likelihood maps presented here were generated by Resnet101 networks that were trained and tested on retinotopic space using the log-polar grid. (A) The original image, it is processed on a single iteration, with the likelihood maps displayed in accordance with the specific combinations of labels of interest. (B) Likelihood maps indicating the probability of the presence of the labels “dog”, “cat” and “bird” in the image. (C) The sample derived from all fixation points that yield the highest likelihood value for a label of interest.
Multi-label and multi-task extension
@@
The likelihood map can be extended along the dimension of labels. The networks perform a discriminative categorization among $1000$ ImageNet labels, and the likelihood maps generated by the model vary depending on the label or class being predicted for a given input image. Specifically, the spatial distribution of high likelihood regions differs according to the object category the model is trying to identify. This suggests that the model learns to focus on different discriminative regions for different classes, providing insight into how the model’s spatial attention adapts to the visual patterns that drive each classification decision. We tested this hypothesis on an example image by generating likelihood maps for the classes “dog”, “cat”, and “bird”. As shown in Figure 10, the peak activations in these maps appear in distinct spatial regions corresponding to the locations of each respective animal in the image.
@@ Discussion
The first and main result of this study is to demonstrate the excellent ability of standard deep CNNs to deal with foveated retinotopic inputs, even though this mapping enforces a radical transformation of the visual inputs. ResNet networks easily adapt to these inputs and the accuracy rates achieved with retinotopic inputs are equivalent to those of the original models. This is surprising given that the networks used in this re-training process were previously trained on Cartesian images. Images with a log-polar transformation show a high degree of distortion, particularly a high compression of visual information around the fixation point and a degradation of textures in the periphery (see Figure 1). One possible hypothesis is that the degradation of texture during the frame of reference change may cause the network to rely on shape rather than texture (da Costa et al., 2024). A further study could entail a comparison of these visual processing parameters with those of human vision, thereby providing insight into their evolutionary trajectory.
@@
In addition, the log-polar transformation has the advantage of better invariance to zooms and rotations. First, this study shows that for the original ResNet network on “regular” images, the average accuracy with a zoom or a rotation dropped sharply compared to baseline accuracy, confirming that these simple geometrical transformations misled the networks. For rotations, the decline was steepest around $160^{\circ}$, demonstrating limited rotational invariance compared to humans (Rousselet et al., 2003; Guyonneau et al., 2006). The robustness differs slightly when studying the impact of zooms. This may be attributed to the transformation under examination, which already provides a zooming-in effect of the fovea that accentuates the information in the area surrounding the point of fixation. To test this hypothesis, we could train a retinotopic version without applying the logarithmic function to the eccentricity axis. Finally, this invariance to rotations and zooms comes at the cost of reduced invariance to translations. For images not centered on the region of interest, one would need to shift the fixation point to the region of interest, similar to eye saccades in biological vision. The effects of this trade-off can be observed in the comparison between the “regular” version of the networks and the “focus” version.
In addition, the log–polar transformation confers improved invariance to zoom and rotation. In the original ResNet trained on “regular” Cartesian images, accuracy declined sharply under zooms or rotations, confirming that even simple geometric transformations can mislead the network. The steepest drop occurred around $160^{\circ}$, highlighting limited rotational invariance compared to humans (Rousselet et al., 2003; Guyonneau et al., 2006). Robustness to zoom was less affected, likely because the transformation itself magnifies foveal input relative to the periphery. This robustness, however, comes at the cost of reduced tolerance to translations: images misaligned with the fixation point require shifting the sampling center, analogous to saccades in biological vision. This trade-off is evident when comparing the “regular” and “focus” network variants. Similar effects have been reported previously:  Pun and Lee showed that log–polar wavelet signatures enhance joint scale–rotation invariance, while  Amorim et al. demonstrated improved rotational robustness in natural image categorization and highlighted the positional sensitivity introduced by inhomogeneous log–polar sampling. Finally, previous research has demonstrated that improved invariance to scale variations is often associated with greater robustness to input compression (Remmelzwaal et al., 2019). This dual robustness — the ability to maintain performance across scaled inputs and compressed representations — facilitates the application of categorisation models to high-resolution images without compromising accuracy. These properties are particularly valuable in computational neuroscience, where visual processing systems must generalise across diverse input conditions.
@@
The second result of this study is the emergence of localization properties in networks re-trained with foveated inputs. Through the definition of likelihood maps, which collate the model’s output while scanning the visual scene at a limited number of fixation points, we gain insight into the specificities of retinotopic processing. This transformation provides a more focused view, thus better separating different elements of the image when focusing on specific parts. Such processing is reminiscent of pre-attentive mechanisms which allow biological vision to selectively process important zones of the visual space. As a result, retinotopic transformation provides a proxy for the measurement of saliency, particularly with respect to a set of cued labels.
@@
One hypothesis is that the foveated retinotopic mapping implemented in the log-polar transform provides an efficient prior for visual object geometry. Indeed, a main source of variability in the view of an object comes from displacements of the observer relative to it, for example as the object scales in the visual field as one approaches it, or as the object rotates with a rotation of the head. The log-polar mapping allows for a more invariant representation which explicitly implements this prior for these displacements. Additionally, imposing translation invariance of the representation establishes a prior on the possible representations allowed by the network (Lempitsky et al., 2018). However, it has been observed that different classes have different statistics, and the relative size of buses is on average larger than that of cats. Consequently, we expect that different foveated retinotopies may emerge in different ecological niches.
@@
In our case, it seems that the foveated retinotopic mapping allows for a more precise localization of the category of interest compared to off-the-shelf pre-trained networks using a Cartesian representation. It also gives us insight into the features on which our networks actually rely. Such information can be compared with physiological data (Crouzet, 2011), used to design better CNNs, and ultimately enable the development of physiological tests to further explore the features needed to classify a label of interest. In particular, by focusing on the point of fixation with the highest probability in the likelihood maps instead of the center of the bounding box defined in the “focus” dataset, we could consider refining the training of the network using our retinotopic mapping in a semi-supervised fashion.
Perspectives and future work
@@
Building on these observations, simulating human-like saccadic eye movements offers a powerful approach to probe network mechanisms. A protocol that iteratively classifies foveated image patches could approximate natural viewing and reveal how performance varies across the visual field, as hypothesized in psychophysical experiments (Han et al., ). Comparing accuracy under different saccadic planning strategies—for example, prioritizing the most uncertain versus the most likely locations—would shed light on network attentional dynamics. This framework also enables modeling of biologically inspired saccade generation, allowing direct comparison with human visual search behavior.
@@
Overall, implementing foveated classification with algorithmic saccades would provide a powerful method for validating existing attentional mechanisms in these networks, as well as inspiring new architectural innovations through embodied, task-driven visual attention modeling. In particular, this line of research would be especially suitable for dual-pathway models that effectively infer ego-motions (Bakhtiari et al., 2021; Mineault et al., 2021). Finally, the implementation of this robust categorization, coupled with refined localization of a label of interest and optimal saccade selection, could allow us to extend this study to more complex tasks. Visual search (i.e., the simultaneous localization and detection of a visual target) represents one such application where likelihood maps could provide the underlying pre-attentive mechanisms on which its effectiveness depends.
@@
In most mammals and amphibians, the arrangement of the (external) visual field is preserved in the early visual pathway, a feature called retinotopy. Retinotopic mapping results from the combined effect of the arrangement of photoreceptors in the retina and their output convergence via the optic nerve. This causes nearby regions of the visual field to activate adjacent neural structures as signals travel from the retina to the brain. These mappings differ from species to species, and our study concentrates on foveated vision (as in humans) which gives more resolution to the central field of view. In particular, we implement it by transforming the Cartesian pixel coordinates into log-polar coordinates (Javier Traver and Bernardino, 2010).
This simple parameterized transformation models accurately this biologically inspired retinotopic mapping. Considering arbitrary images (potentially with multiple channels such as RGB), each pixel’s position is defined by two real coordinates $(x,y)$ on a Cartesian reference frame. By convention, $x$ and $y$ are here considered belonging to the interval [1,1][-1,1], with $(0,0)$ being the center of the image. To implement the concentration of pixels near the center of the retina, we need to consider an irregular grid in the Cartesian referential that maps to a regular grid in the log-polar referential. In the log-polar referential, the location of each pixel has corresponding coordinates (${\rho},{\theta}$) as defined by previous studies (Hao et al., 2021) by:
@@
with $(x_{0},y_{0})$ defining the “center of fixation” (see Figure 1(Traver and Pla, 2003). Importantly, $\rho$ and $\theta$ are only defined for $(x,y)\neq(x_{0},y_{0})$. In most of our experiments, we consider $x_{0}=(0,0)$, allowing us to focus on the central part of the image. Each dot ($\rho$, $\theta$) in the log-polar coordinate system has thus a unique correspondence in the Cartesian coordinate system (and vice versa): For each ($\log\rho$, $\theta$) belonging to the grid, the corresponding pixel coordinate is $(x_{0}+\rho\cos\theta,y_{0}+\rho\sin\theta)$.
In practice, images have a finite resolution and to avoid biases in the evaluation between networks, the number of angles sampled ($N_{\theta}$) and the number of eccentricities sampled ($N_{\rho}$) are set to $224$, so that the size of the transformed image match the resolution $224\times 224$ of the input images. Note that native image resolution is generally higher than that used during network processing, with the average resolution in the Imagenet dataset being around $500\times 500$ pixels. This transformation is performed with the PyTorch library (Paszke et al., 2019) through the use of the grid_sample() function, which maps the pixels of an input image to the coordinates of any arbitrary grid, using a linear interpolation to estimate the value of the pixels. This function is used, for instance, in spatial transformer networks (Jaderberg et al., 2016).
Let’s now define each coordinates. All $\theta$ values are within a linear distribution in $0;2\pi)$, while $\rho$ values are within a logarithmic interval with $r_{\text{min}}$ the minimal radius and $r_{\text{max}}$ be the maximal radius (with $N_{\rho}$ the radial resolution). In practice, we use a log-polar grid with an outer log-radius of $\log_{2}r_{\text{max}}=0$ ($r_{\text{max}}=1$, defining a circle tangent to the image box) and an inner log-radius of $\log_{2}r_{\text{min}}=-5$ ($r_{\text{min}}=2^{-5}$). In summary, the regular grid is the interval $\log_{2}r_{\text{min}},...,\log_{2}(r_{\text{min}}+i\times\frac{r_{\text{max}}-r_{\text{min}}}{N_{\rho}-1}),...,\log_{2}r_{\text{max}}]$, for $i$ in $0,...,N_{\rho}-1]$, in the $\log_{2}\rho$ dimension and $0,...,j\times\frac{2\pi}{N_{\theta}},...,2\pi(1-\frac{1}{N_{\theta}})]$, for $j$ in $0,...,N_{\theta}-1]$ in the $\theta$ dimension. Note that some pixels in the log-polar grid may be smaller than the pixels from the Cartesian grid.
@@
Convolutional neural networks (CNNs) have become essential tools in image classification, with several pre-trained models available for download. For example, the VGG family, including VGG16 and VGG19, introduced by Simonyan and Zisserman (2015), uses deep architectures with $16$ or $19$ layers, consisting of stacked convolutional layers followed by fully connected layers. ResNet (Residual Networks), introduced by He et al. (2015), addresses the vanishing gradient problem in deep networks by incorporating skip connections, allowing for the training of extremely deep networks (e.g. ResNet $50$, ResNet $101$ with respectively $50$ and $101$ layers). These models are widely used due to their robustness and scalability. The implementation of these deeper networks have demonstrated that deeper networks display enhanced resilience; however, this improvement is coupled with an overall increase in computational complexity (He et al., 2015). Therefore, based on these findings, we focus on the deep CNN ResNet (with $18$ to $101$ layers from the PyTorch library  [^1] We further introduced a circular padding in the convolutions, however we controlled this had little impact overall.) on the ImageNet-1K (Russakovsky et al., 2015) categorization challenge which consists in classifying natural images into $1000$ labels.
@@
Figure 11: (A) Original image from the ImageNet dataset. (B) Bounding-box annotation shown as a heat map. (C) Downsampled heat map ($11\times 11$) used as ground truth for localization. The “regular” dataset uses the full image as input (Cartesian or log-polar for the retinotopic framework), whereas the “focus” dataset uses the cropped bounding box of the target label (yellow boxes in B). To ensure consistency, if there are multiple boxes present, only the first annotation provided in the dataset is used when constructing the “focus” dataset.
@@
Two datasets were used for our study: the first dataset is the one from the ImageNet (Russakovsky et al., 2015) challenge, which is the most widely benchmarked due to its huge collection of images and associated labels (the subset of ImageNet used in this study, i.e. ImageNet-1K with $1000$ labels and about $1000$ examples per label). In addition to the classification task, we consider here the localization task, which is prominent in computer vision. It consists in identifying both the label and the position of one or more objects of interest in an image. A distinction can be made between (i) the visual search task, where the label is given in advance and the task is simply to find the position of the object in the image, and (ii) the image labeling task, which consists in identifying both the objects and their position in the image, in the form of a labeled bounding box (see Figure 11). Despite its advantages, ImageNet has limitations for localization tasks. For example, the dataset lacks multi-labeling, with only one label per image. It is worth mentioning that ImageNet has some biases, the objects of interest are generally centered in the images, and the proportion of bounding boxes relative to the image size is relatively large, which may limit the impact of certain analyses.